OpenStudy (anonymous):

Write the equation of the line having slope=3/4, and passing through (-4,-2). GRAPH the line.

OpenStudy (anonymous):

y=(3/4)x+b -2=(3/4*-4)+b -2=-3+b b=1 y=(3/4)x+1